A sample of hydrogen gas occupies a volume of 9.64 L at 58.0°C and 312 torr. If the volume of the gas sample is decreased to 8.05 L, while its temperature is decreased to –9.0°C, the resulting gas pressure will be torr.
Answer #1

            Initial                                                                          Final

          V1 = 9.64L                                                                 V2 = 8.05L

            P1 = 312torr                                                             P2 =

           T1   = 58+273   = 331K                                               T2 = -9+273   = 264K

                         P1V1/T1    =     P2V2/T2

                              P2           =    P1V1T2/T1V2

                                               =   312*9.64*264/(331*8.05)

                                                  = 298 torr

              The pressure of gas = 298 torr >>>>>answer